# ToMCCA-3: A realistic 3-body coalescence model

### v1

@@ Abstract
The formation of light nuclei in high-energy collisions provides valuable insights into the underlying dynamics of the strong interaction and the structure of the particle-emitting source. Understanding this process is crucial not only for nuclear physics but also for astrophysical studies, where the production of rare antinuclei could serve as a probe for new physics. This work presents a three-body coalescence model based on the Wigner function formalism, offering a refined description of light-nucleus production. By incorporating realistic two- and three-body nuclear interaction potentials constrained by modern scattering and femtoscopic correlation data, our approach improves on traditional coalescence models. The framework is validated using event generators applied to proton-proton collisions at $\sqrt{s}=13$ TeV to predict the momentum spectra of light (anti) nuclear nuclei with mass number $A=3$, which are then compared with the experimental data from ALICE. Our results demonstrate the sensitivity of light nucleus yields to the choice of nuclear wave functions, emphasizing the importance of an accurate description of the coalescence process. This model lays the foundation for the extension of coalescence studies of $A=3$ light nuclei to a wider range of collision systems and energies.
[…]
subalign[1][]
[…]
spliteq
$\BODY\BODY\begin{split}\BODY\end{split}$ (2)
@@ 1 Introduction
This question has received particular attention at high energy hadron colliders such as the Large Hadron Collider (LHC) and the Relativistic Heavy Ion Collider (RHIC). Due to the large integrated luminosities delivered over the past decade, multi-differential measurements of light nuclei and antinuclei production could be performed with unprecedented precision. New techniques, such as femtoscopy [1, 2] and correlation studies, have provided deeper insights into the production mechanisms of these bound states.
Light (anti-)nuclei yields and particle ratios measured as functions of the final-state charged-particle multiplicity pseudorapidity density (or multiplicity) have evidenced the dependence of nucleus production on the size of the particle emitting source [3, 4, 5, 6, 7]. Different particle species, among which nucleons and strange baryons, are observed to be emitted from a common particle source in proton-proton collisions at center of mass energies of $\sqrt{s}=13$ TeV [8, 9], whose radius decreases with transverse momentum [^1] More precisely, the one dimensional radius assuming a gaussian source profile decreases with the average particle pair transverse momentum, $k_{T}$.. In addition, pion kaon and proton radii are observed to increase with multiplicity from pp to p–Pb (small systems) to heavy-ion collisions (large systems[10, 11]. Moreover, proton-proton (p–p), proton-deuteron (p–d), Lambda-deuteron ($\Lambda$–d) and three particle correlations (p–p–p and p–p–$\Lambda$) studied with femtoscopy have allowed access to the strong interaction between baryons, shedding light on the three-body forces acting in the system of two-and three-nucleons [12].
@@
Among the light nuclei and antinuclei, deuterons have been measured with the highest precision, as a function of multiple variables such as multiplicity, momentum, and rapidity. Models attempting to describe their formation—including statistical hadronization [13, 14, 15] and coalescence [16, 17, 18, 19, 20], a nuclear fusion-like process in high-energy collisions—have been largely successful, despite relying on different underlying principles. In statistical hadronization models (SHM), light nuclei are produced from a hadron gas in thermal equilibrium and according to the laws of quantum statistics and with an appropriate treatment of the quantum number conservation laws depending on the size of the system. SHM predicted abundances depend on the particle mass, spin, temperature and volume of the system. The coalescence approach describes how individual (anti)nucleons merge into (anti)nuclei based on their proximity in momentum and spatial coordinates. However, discrepancies between the measured and predicted observables remain, particularly in the production of heavier systems such as helium, tritium, and hypertriton (a bound state of a neutron, a proton, and a $\Lambda$ baryon) [7, 21]. The latter confirms to be a key observable for discriminating between thermal and coalescence production [22, 23, 24] due to its broad size ($\sim 10$ fm) and small energy ($\sim 150$ keV) that binds the $\Lambda$ to the deuteron core [25]. The first hypertriton measurement in p–Pb collisions, supports the coalescence model and disfavors the hypothesis of thermal production [21].
While widely used, coalescence models often involve free parameters—such as the coalescence momentum—whose values are not derived from first principles and vary depending on the production environment. Traditional models assume simple momentum-based coalescence, while more advanced quantum-mechanical approaches incorporate nucleon wave functions and the size of the emission source, according to a Wigner function-based formulation [19, 20, 24]. Such an approach also finds application in recent developments based on Monte Carlo methods that allow us to apply coalescence as an afterburner to particle production [26, 27].
@@
One of the most intriguing applications of the research on (anti)nuclei formation mechanisms is the indirect detection of dark matter (DM) through cosmic antinuclei, such as antideuterons and antihelions. Unlike standard cosmic ray antimatter (e.g., antiprotons), these antinuclei have a significantly suppressed background from astrophysical sources, making them a promising signature of DM annihilation [28, 29, 30, 31, 32, 33]. However, their production mechanisms remain uncertain, and data from particle accelerators become a key input. The tentative measurements of ${}^{3,4}\overline{\mathrm{He}}$ antinuclei in cosmic rays by the AMS-02 Collaboration have sparked several studies in order to interpret these findings. Although the experimental results have not been published in a peer-reviewed journal yet, the confirmation of even one such antinucleus would challenge the current understanding of cosmology [34, 35].
@@
In this paper, the coalescence model ToMCCA [36] is extended to the $A=3$ case. The predictions include not only the nuclei 3He and 3H, but also the hypernucleus ${}^{3}_{\Lambda}$. The Wigner function formalism has been implemented for $A=3$ nuclei using realistic wave functions based on the pair-correlated hyperspherical harmonics (PHH) method [37]. For 2-body nucleon-nucleon (NN) interactions, the Argonne $\text{Argonne }v_{18}$ potential (AV18) [38] is used, together with the Urbana IX (UIX) 3-body interaction [39]. Furthermore, the Minnesota potential [40] is tested as a simple 2-body potential which is able to reproduce the binding energy 3He and 3H. For ${}^{3}_{\Lambda}$, two approaches are considered: The Congleton [41] approach is a simplified model of the ${}^{3}_{\Lambda}$, with a deuteron at its core, modeled using the Argonne $\text{Argonne }v_{18}$ wave function of the deuteron, and a shallowly bound $\Lambda$. The parameters of this model were tuned to reproduce modern form factor calculations [24, 42]. The second approach is based on central Gaussian interactions, including 2 and 3-body potentials [43]. The parameters of the 2-body potentials are tuned to reproduce experimental scattering lengths and effective ranges in the NN case as well as the ones from $\chi$EFT in the N-$\Lambda$ case. The parameters of the 3-body potential are chosen to match a binding energy of $B({}^{3}_{\Lambda}\text{H})=2.40$ MeV. The ${}^{3}_{\Lambda}$ wave function is calculated employing the hyperspherical harmonic method [37, 44]. Details of this calculation are discussed in this paper.
@@ 2 Coalescence approach for A = 3 nuclei
The momentum spectra of light (anti)(hyper)nuclei [^2] The formation mechanism is assumed to be the same for nuclei and their antimatter counterpart, as well as for hypernuclei. In the following, the (anti) and (hyper) prefixes will be omitted. with mass number $A=3$, such as 3H, 3He, and ${}^{3}_{\Lambda}$, can be computed using the coalescence formalism. In this framework, light nuclei are assumed to form in high-energy hadronic collisions through the coalescence of nucleons and hyperons. A key assumption is that interactions between coalescing hadrons and non-participating hadrons are subdominant. We develop a coalescence model for $A=3$ light nuclei based on the state-of-the-art Wigner function formalism [26, 27]. We compute Lorentz-invariant momentum spectra of 3H, 3He, and ${}^{3}_{\Lambda}$ using the three-particle density matrix formalism as follows: [^3] Note that in these calculations, we use bold and italic fonts for three- and four-vectors.
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\leavevmode\nobreak\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{4}}\int\mathrm{\leavevmode\nobreak\ d}^{4}x_{1}\int\mathrm{\leavevmode\nobreak\ d}^{4}x_{2},\int\mathrm{\leavevmode\nobreak\ d}^{4}x_{3}\int\mathrm{\leavevmode\nobreak\ d}^{4}x_{1}^{\prime}\int\mathrm{d}^{4}x_{2}^{\prime}\int\mathrm{d}^{4}x_{3}^{\prime}\\ \times\Psi^{*}\left(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime},\right)\Psi\left(x_{1},x_{2},x_{3}\right)\rho_{1,2,3}\left(x_{1},x_{2},x_{3};x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime}\right),$ (3)
where $\Psi\left(x_{1},x_{2},x_{3}\right)$ is the three-particle bound state Bethe Salpeter amplitude and $\rho_{1,2,3}$ is the three-particle reduced density matrix, while $S_{\mathrm{A}}$ accounts for the spin and isospin statistics for all three cases. The density matrix for three particle systems is assumed to be factored into single particle densities [24], $\rho_{1,2,3}\left(x_{1},x_{2},x_{3};x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime}\right)=\rho_{1}\left(x_{1};x_{1}^{\prime}\right)\times\rho_{1}\left(x_{2};x_{2}^{\prime}\right)\times\rho_{1}\left(x_{3};x_{3}^{\prime}\right)$, and the single-particle density is written in terms of the single particle Wigner function, $f_{1}^{W}$. One can make the assumption
@@
At this stage, making both the low-energy approximation and the equal-time approximation (see App. A), Eq. (3) takes the form of a yield equation for $A=3$ cluster in three-dimensional space. The yield in three-dimensional space is written as
$\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\leavevmode\nobreak\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{12}}\int\mathrm{d}^{3}\bm{p}_{1}\int\mathrm{d}^{3}\bm{p}_{2}\int\mathrm{d}^{3}\bm{p}_{3}\int\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}_{1}\int\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}_{2}\,\int\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}_{3}\int\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}_{1}^{\prime}\int\mathrm{d}^{3}\bm{x}_{2}^{\prime}\int\mathrm{d}^{3}\bm{x}_{3}^{\prime}\\ \times\Psi^{*}_{NR}\left(\bm{x}_{1}^{\prime},\bm{x}_{2}^{\prime},\bm{x}_{3}^{\prime},\right)\Psi_{NR}\left(\bm{x}_{1},\bm{x}_{2},\bm{x}_{3}\right)\,e^{-i\bm{p}_{1}\cdot\left(\bm{x}_{1}-\bm{x}_{1}^{\prime}\right)-i\bm{p}_{2}\cdot\left(\bm{x}_{2}-\bm{x}_{2}^{\prime}\right)-i\bm{p}_{3}\cdot\left(\bm{x}_{3}-\bm{x}_{3}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3},\frac{\bm{x}_{1}+\bm{x}_{1}^{\prime}}{2},\frac{\bm{x}_{2}+\bm{x}_{2}^{\prime}}{2},\frac{\bm{x}_{3}+\bm{x}_{3}^{\prime}}{2}\right)\,,$ (6)
@@
Let us first consider a system consisting of two identical particles with masses $m_{1}=m_{2}=M$ and a third particle with a different mass $m_{3}$. We define the mass ratio as $\kappa=\frac{m_{3}}{M}$ so that the cases of 3He and 3H can be simply recovered by setting $\kappa=1$ and the case of ${}^{3}_{\Lambda}$ by setting $\kappa=\frac{m_{\Lambda}}{M}\approx 1.188$ in the final expressions. The Jacobi vectors are
@@
${\bm{P}}={\bm{p}}_{1}+{\bm{p}}_{2}+{\bm{p}}_{3}\ ,\qquad\bm{k}_{2}={1\over 2}\Bigl{[}{\bm{p}}_{2}-{\bm{p}}_{1}\Bigr{]}\ ,\qquad\bm{k}_{1}=\sqrt{\kappa\over 2+\kappa}\left[{{\bm{p}}_{3}\over\kappa}-{{\bm{p}}_{1}+{\bm{p}}_{2}\over 2}\right]\ .$ (11)
@@
It is easy to show that
@@
where function $H_{\mathrm{A}}$ is assumed to be the product of Gaussian distributions $H_{\mathrm{A}}=\prod_{i=1}^{\mathrm{A}}h(x_{i})$ and $h(x_{i})$ is a normalized Gaussian distribution,
@@
The width parameters $\sigma$ is related to the particle emission profile and can be obtained from two-particle femtoscopic source studies in ultra-high-energy hadronic collisions [9, 45, 46]. The function $G_{\mathrm{A}}$ is still not known theoretically; however, since it is a distribution related to the momentum spectra of the individual nucleons, the values of this function can be incorporated by employing the event generators [26].
@@ 2.1 Generalized 3-body coalescence approach
In the following, we will derive the calculation of the coalescence probability for ${}^{3}_{\Lambda}$. This defines a generalized approach for 2 equal (nucleons) and one different (hyperon) particle. Under the assumption that the third particle is equal to the other two, the cases of 3He and 3H are derived. We consider two distinct approaches for hypertriton wave function calculations. The first employs the state-of-the-art hyperspherical harmonics (HH) approach [37, 44] to solve the three-body bound-state Schrödinger equation, utilizing model Gaussian p–$\Lambda$ potentials devised in Ref. [43]. These potentials are constructed such to reproduce the low-energy scattering data as obtained from realistic hyperon-nucleon interactions devised within a chiral effective field theory framework [47, 48, 49]. These latter interactions are constrained by the available N–$\Lambda$ scattering data [50] and the latest femtoscopic measurements of the p–$\Lambda$ correlation function [50]. Although rather simple, with the inclusion of N–N–$\Lambda$ three-body potentials chosen to reproduce the hypertriton binding energy, these Gaussian potentials well decribe both p–$\Lambda$ and p–p–$\Lambda$ correlation functions [43]. In addition to them, we have also considered the N–$\Lambda$ and N–N-$\Lambda$ potentials devised in Ref. [51, 52]. The precise form of the potentials and the HH method to compute the wave functions is dicussed in detail in Appendix B. The second approach is based on the work developed by Congleton et al. [41], as motivated in Ref. [53].
The Wigner density matrix for the system of $\mathrm{np}\Lambda$ appearing in Eq. (53) can be written as
$W_{\mathrm{{np}\Lambda}}\left(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3},\frac{{\bm{x}}_{1}+{\bm{x}}_{1}^{\prime}}{2},\frac{{\bm{x}}_{2}+{\bm{x}}_{2}^{\prime}}{2},\frac{{\bm{x}}_{3}+{\bm{x}}_{3}^{\prime}}{2}\right)$ $=$ ${1\over(2\pi\sigma^{2})^{3}}e^{-{({\bm{x}}_{1}+{\bm{x}}_{1}^{\prime})^{2}+({\bm{x}}_{2}+{\bm{x}}_{2}^{\prime})^{2}\over 8\sigma^{2}}}$ (19)
×$\times{1\over(2\pi\sigma^{\prime 2})^{3/2}}e^{-{({\bm{x}}_{3}+{\bm{x}}_{3}^{\prime})^{2}\over 8\sigma^{\prime 2}}}G_{\mathrm{np}\Lambda}({\bm{p}}_{1},\bm{p}_{2},{\bm{p}}_{3})\ ,$
where $\sigma$ and $\sigma^{\prime}$ represent the femtoscopic source sizes for N–N pairs and N–$\Lambda$ pairs, respectively, and are assumed to be independent of each other. The values of $\sigma$ and $\sigma^{\prime}$ are constrained by recent source size measurements and can be obtained from event generators, which are tuned to reproduce the experimental source size measurements; more details are discussed in the subsequent sections.
In the following, we define $\tau=(\sigma/\sigma^{\prime})^{2}$. With this definition, one obtains
$e^{-{({\bm{x}}_{1}+{\bm{x}}_{1}^{\prime})^{2}+(\bm{x}_{2}+\bm{x}_{2}^{\prime})^{2}\over 8\sigma^{2}}}e^{-{(\bm{x}_{3}+\bm{x}_{3}^{\prime})^{2}\over 8\sigma^{\prime 2}}}=e^{-{(\bm{x}_{1}+\bm{x}_{1}^{\prime})^{2}+(\bm{x}_{2}+\bm{x}_{2}^{\prime})^{2}+\tau\,(\bm{x}_{3}+\bm{x}_{3}^{\prime})^{2}\over 8\sigma^{2}}}\ .$ (20)

### v2

@@ Abstract
The formation of light (anti)nuclei in high-energy collisions provides valuable insights into the underlying dynamics of the strong interaction and the structure of the particle-emitting source. Understanding this process is crucial not only for nuclear physics but also for astrophysical studies, where the production of rare antinuclei could serve as a probe for new physics. This work presents a three-body coalescence model based on the Wigner function formalism, offering a refined description of light-(anti)nuclei production in ultra relativistic proton-proton collisions. By incorporating realistic two- and three-body nuclear interaction potentials constrained by modern scattering and femtoscopic correlation data, our approach improves on traditional coalescence models. The framework is validated using the ToMCCA event generator applied to proton-proton collisions at $\sqrt{s}=13$ TeV to predict the momentum spectra of light (anti)(hyper)nuclei with mass number $A=3$, which are then compared with the experimental data from ALICE. Our results demonstrate the sensitivity of light nuclei yields to the choice of nuclear wave functions, emphasizing the importance of an accurate description of the coalescence process. This model lays the foundation for the extension of coalescence studies of $A=3$ light nuclei to a wider range of collision systems and energies.
@@ 1 Introduction
This question has received particular attention at high energy hadron colliders such as the Large Hadron Collider (LHC) and the Relativistic Heavy Ion Collider (RHIC). Due to the large integrated luminosities delivered over the past decade, multi-differential measurements of light nuclei and antinuclei production could be performed with unprecedented precision. New techniques, such as femtoscopy [40, 33] and correlation studies [17, 20], have provided deeper insights into the production mechanisms of these bound states.
[…]
@@
Among the light nuclei and antinuclei, deuterons have been measured with the highest precision, as a function of multiple variables such as multiplicity, transverse momentum, and rapidity. Models attempting to describe their formation—including statistical hadronization [21, 22, 51] and coalescence [27, 31, 44, 46, 26], a nuclear fusion-like process in high-energy collisions – have been largely successful, despite relying on different underlying principles. In statistical hadronization models (SHM), light nuclei are produced from a hadron gas in thermal equilibrium, according to the laws of quantum statistics, and with an appropriate treatment of the quantum number conservation laws, depending on the size of the system. SHM-predicted abundances depend on the particle mass and spin, as well as the temperature and volume of the system. The coalescence approach describes how individual (anti)nucleons merge into (anti)nuclei based on their proximity in momentum and spatial coordinates. However, discrepancies between the measured and predicted observables remain, particularly in the production of heavier systems such as helium, triton, and hypertriton (a bound state of a neutron, a proton, and a $\Lambda$ baryon) [11, 7]. The latter confirms to be a key observable for discriminating between thermal and coalescence production [47, 24, 23] due to its broad size ($\sim 10$ fm) and small energy that binds the $\Lambda$ to the deuteron core (world average $B_{\Lambda}=105\pm 26$ keV, [29]). The first hypertriton measurement in p–Pb collisions, supports the coalescence model and disfavors the hypothesis of thermal production [7].
@@
While widely used, coalescence models often involve free parameters, such as the coalescence momentum $p_{0}$, whose values are not derived from first principles and vary depending on the production environment. Traditional models assume simple momentum-based coalescence, while more advanced quantum-mechanical approaches incorporate nucleus wave functions and the size of the emission source, according to a Wigner function-based formulation [46, 26, 23]. Such an approach also finds application in recent developments based on Monte Carlo methods that allow us to apply coalescence as an afterburner to particle production [37, 41]. One of the most intriguing applications of the research on (anti)nuclei formation mechanisms is the indirect detection of dark matter (DM) through cosmic antinuclei, such as antideuterons and antihelium. Unlike standard cosmic ray antimatter (antiprotons and positrons), these antinuclei have a significantly suppressed background from astrophysical sources, making them a promising signature of DM annihilation [32, 36, 30, 28, 39, 25]. However, their production mechanisms remain uncertain, and measurements from particle accelerators become a key input. The tentative measurements of $^{3,4}\overline{\mathrm{He}}$ antinuclei in cosmic rays by the AMS-02 Collaboration have sparked several studies in order to interpret these findings. Although the experimental results have not been published in a peer-reviewed journal yet, the confirmation of even one such antinucleus would challenge the current understanding of cosmology [50, 12].
In this paper, the coalescence model ToMCCA [42] is extended to the $A=3$ case. ToMCCA is a lightweight event-by-event coalescence Monte-Carlo generator designed to study light-(anti)nuclei formation without the overhead of a full event generator and with easy customization. It includes parameterizations of the multiplicity dependence of proton transverse-momentum spectra and of the baryon-emitting source size tuned to (anti)deuteron measurements by ALICE in pp collisions at $\sqrt{s}=5-13\,\mathrm{TeV}$, as well as event-multiplicity distributions taken from EPOS. ToMCCA reproduces all the deuteron $p_{\mathrm{T}}$ distributions by ALICE within $\sim 5\%$.
@@
The predictions include not only the nuclei 3He and 3H, but also the hypernucleus $^{3}_{\Lambda}$H. The Wigner function formalism has been implemented for $A=3$ (anti)nuclei using realistic wave functions based on the pair-correlated hyperspherical harmonics (PHH) method [38]. For 2-body nucleon-nucleon (NN) interactions, the $\text{Argonne }v_{18}$ potential (AV18) [52] is used, together with the Urbana IX (UIX) 3-body interaction [45]. Furthermore, the Minnesota potential [49] is tested as a simple 2-body potential which is able to reproduce the binding energies of 3He and 3H. For $^{3}_{\Lambda}$H, a simplified approach is considered: the Congleton [34] approach is a simplified model of the $^{3}_{\Lambda}$H, with a deuteron at its core, modeled using the $\text{Argonne }v_{18}$ wave function of the deuteron, and a shallowly bound $\Lambda$. The parameters of this model were tuned to reproduce modern form factor calculations [23, 35], and it uses a $\Lambda$ binding energy of $B_{\Lambda}=128$ keV, well in line with the world average.
@@ 2 Coalescence approach for A = 3 nuclei
The momentum spectra of light (anti)(hyper)nuclei [^2] The formation mechanism is assumed to be the same for nuclei and their antimatter counterpart, as well as for hypernuclei. In the following, the (anti) and (hyper) prefixes will be omitted. with mass number $A=3$, such as 3H, 3He, and $^{3}_{\Lambda}$H, can be computed using the coalescence formalism. In this framework, light nuclei are assumed to form in high-energy hadronic collisions through the coalescence of nucleons and hyperons. A key assumption is that interactions between coalescing hadrons and non-participating hadrons are subdominant. We develop a coalescence model for $A=3$ light nuclei based on the Wigner function formalism [37, 41]. We compute Lorentz-invariant momentum spectra of 3H, 3He, and $^{3}_{\Lambda}$H using the three-particle density matrix formalism as follows: [^3] Note that in these calculations, we use bold and italic fonts for three- and four-vectors.
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\penalty 10000\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{4}}\int\mathrm{\penalty 10000\ d}^{4}x_{1}\int\mathrm{\penalty 10000\ d}^{4}x_{2},\int\mathrm{\penalty 10000\ d}^{4}x_{3}\int\mathrm{\penalty 10000\ d}^{4}x_{1}^{\prime}\int\mathrm{d}^{4}x_{2}^{\prime}\int\mathrm{d}^{4}x_{3}^{\prime}\\ \times\Psi^{*}\left(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime},\right)\Psi\left(x_{1},x_{2},x_{3}\right)\rho_{1,2,3}\left(x_{1},x_{2},x_{3};x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime}\right),$ (1)
[…]
@@
At this stage, making both the low-energy approximation and the equal-time approximation (see App. E), Eq. (1) takes the form of a yield equation for $A=3$ cluster in three-dimensional space. The yield in three-dimensional space is written as
$\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\penalty 10000\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{12}}\int\mathrm{d}^{3}\bm{p}_{1}\int\mathrm{d}^{3}\bm{p}_{2}\int\mathrm{d}^{3}\bm{p}_{3}\int\mathrm{\penalty 10000\ d}^{3}\bm{x}_{1}\int\mathrm{\penalty 10000\ d}^{3}\bm{x}_{2}\,\int\mathrm{\penalty 10000\ d}^{3}\bm{x}_{3}\int\mathrm{\penalty 10000\ d}^{3}\bm{x}_{1}^{\prime}\int\mathrm{d}^{3}\bm{x}_{2}^{\prime}\int\mathrm{d}^{3}\bm{x}_{3}^{\prime}\\ \times\Psi^{*}_{NR}\left(\bm{x}_{1}^{\prime},\bm{x}_{2}^{\prime},\bm{x}_{3}^{\prime},\right)\Psi_{NR}\left(\bm{x}_{1},\bm{x}_{2},\bm{x}_{3}\right)\,e^{-i\bm{p}_{1}\cdot\left(\bm{x}_{1}-\bm{x}_{1}^{\prime}\right)-i\bm{p}_{2}\cdot\left(\bm{x}_{2}-\bm{x}_{2}^{\prime}\right)-i\bm{p}_{3}\cdot\left(\bm{x}_{3}-\bm{x}_{3}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3},\frac{\bm{x}_{1}+\bm{x}_{1}^{\prime}}{2},\frac{\bm{x}_{2}+\bm{x}_{2}^{\prime}}{2},\frac{\bm{x}_{3}+\bm{x}_{3}^{\prime}}{2}\right)\,$, (4)
@@
Let us first consider a system consisting of two identical particles with masses $m_{1}=m_{2}=M$ and a third particle with a different mass $m_{3}$. We define the mass ratio as $\kappa=\frac{m_{3}}{M}$ so that the cases of 3He and 3H can be simply recovered by setting $\kappa=1$ and the case of $^{3}_{\Lambda}$H by setting $\kappa=\frac{m_{\Lambda}}{M}\approx 1.188$ in the final expressions. The Jacobi vectors are
@@
${\bm{P}}={\bm{p}}_{1}+{\bm{p}}_{2}+{\bm{p}}_{3}\ ,\qquad\bm{k}_{2}={1\over 2}\Bigl[{\bm{p}}_{2}-{\bm{p}}_{1}\Bigr]\ ,\,\bm{k}_{1}=\sqrt{\kappa\over 2+\kappa}\left[{{\bm{p}}_{3}\over\kappa}-{{\bm{p}}_{1}+{\bm{p}}_{2}\over 2}\right]\ .$ (9)
@@
From the definitions of the Jacobi coordinates, the following relation immediately follows.
@@
[…]
@@
The width parameter $\sigma$ is related to the particle emission profile in the ultra-high-energy hadronic collisions and can be obtained from two-particle femtoscopic source studies in proton-proton collisions [15, 3, 18]. The exact form of the function $G_{\mathrm{A}}$ is still not known theoretically; however, since it is a distribution related to the momentum spectra of the individual nucleons, the values of this function can be incorporated by employing an event generators [37, 41].
@@ 2.1 Generalized 3-body coalescence approach
In the following, we will derive the calculation of the coalescence probability for an arbitrary $A=3$ (anti)(hyper)nucleus. This defines a generalized approach for 2 equal (nucleons) and one different (hyperon) particle. Under the assumption that the third particle is equal to the other two, the cases of 3He and 3H are recovered. The Wigner density matrix for the system of $\mathrm{np}\Lambda$ appearing in Eq. (55) can be written as
$W_{3}\left(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3},\frac{{\bm{x}}_{1}+{\bm{x}}_{1}^{\prime}}{2},\frac{{\bm{x}}_{2}+{\bm{x}}_{2}^{\prime}}{2},\frac{{\bm{x}}_{3}+{\bm{x}}_{3}^{\prime}}{2}\right)$ $=$ ${1\over(2\pi\sigma^{2})^{3}}\exp\left(-{({\bm{x}}_{1}+{\bm{x}}_{1}^{\prime})^{2}+({\bm{x}}_{2}+{\bm{x}}_{2}^{\prime})^{2}\over 8\sigma^{2}}\right)$ (17)
$\times{1\over(2\pi\sigma^{\prime 2})^{3/2}}\exp\left(-{({\bm{x}}_{3}+{\bm{x}}_{3}^{\prime})^{2}\over 8\sigma^{\prime 2}}\right)G_{3}({\bm{p}}_{1},\bm{p}_{2},{\bm{p}}_{3})\ ,$
[…]
$\exp\left(-{({\bm{x}}_{1}+{\bm{x}}_{1}^{\prime})^{2}+(\bm{x}_{2}+\bm{x}_{2}^{\prime})^{2}\over 8\sigma^{2}}-{(\bm{x}_{3}+\bm{x}_{3}^{\prime})^{2}\over 8\sigma^{\prime 2}}\right)=\exp\left(-{(\bm{x}_{1}+\bm{x}_{1}^{\prime})^{2}+(\bm{x}_{2}+\bm{x}_{2}^{\prime})^{2}+\tau\,(\bm{x}_{3}+\bm{x}_{3}^{\prime})^{2}\over 8\sigma^{2}}\right)\ .$ (18)